the length of a rectangle is 12 cm more than the width. if the area of the rectangle is 220 cm^2, find the dimensions of the rectangle.
Answer by ikleyn(52780)

.
L*W = 220


(W+12)*W = 220


W^2 + 12W - 220 = 0


Factorize left side:


(W+22)*(W-10) = 0


The only positive root W = 10 is the solution.


Answer.  The dimensions of the rectangle are 10 (width) and (10+12) = 22 centimeters (length).